Help! Please, Linear Combinations
0.5x + 2.2y = 9
6x + 0.4y = -22

Answer by mananth(16949)

0.5x + 2.2y = 9
6x + 0.4y = -22
multiply both equations by 10
5x+22y=90
60x+4y=-220
5 x + 22 y = 90 .............1
60 x + 4 y = -220 .............2
Eliminate y
multiply (1)by -4
Multiply (2) by 22
-20 x -88 y = -360
1320 x 88 y = -4840
Add the two equations
1300 x = -5200
/ 1300
x = -4
plug value of x in (1)
5 x + 22 y = 90
-20 + 22 y = 90
22 y = 90 + 20
22 y = 110
y = 5